OpenStudy (anonymous):

Wei-Min makes a square quilt that has an area of 100 square feet. It has 25 square blocks, each the same size. What is the length of each side of a block?

OpenStudy (nurali):

100=25x^2 100/25=x^2 4=x^2 x=2

OpenStudy (nurali):

Welcome to Openstudy.